OpenStudy (emilyjones284):

what is S19 for the arithmetic series 2+9+16+23+...?

jimthompson5910 (jim_thompson5910):

S19 means "find the sum of the first 19 terms"

jimthompson5910 (jim_thompson5910):

So you could painstakingly generate the first 19 terms, then add up them, but there's a much easier way

jimthompson5910 (jim_thompson5910):

we'll need the first term and the common difference first term: a1 = ??? common difference: d = ???

OpenStudy (emilyjones284):

a1=2 and d=7

jimthompson5910 (jim_thompson5910):

so the nth term is an = a1 + d(n-1) an = 2 + 7(n-1) an = 2 + 7n - 7 an = 7n - 5

jimthompson5910 (jim_thompson5910):

You use the first term (a1), the nth term (an) and the number of terms (n) to calculate Sn (the sum of the first n terms)

jimthompson5910 (jim_thompson5910):

By this formula Sn = n(a1 + an)/2

OpenStudy (emilyjones284):

wait im confused why do i need the first 5 terms?

jimthompson5910 (jim_thompson5910):

that -5 doesn't mean the first 5 terms

jimthompson5910 (jim_thompson5910):

an = 7n - 5 helps you find any term you want say the 3rd term, so n = 3 an = 7n - 5 a3 = 7*3 - 5 ... plug in n = 3 a3 = 21 - 5 a3 = 16 surely enough, 16 is indeed the 3rd term

OpenStudy (emilyjones284):

ohhhh never mind I misread something sorry

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

use this formula Sn = n(a1 + an)/2 to find S19

jimthompson5910 (jim_thompson5910):

In this case, a1 = 2 an = 7n - 5 (just found this) n = 19

OpenStudy (emilyjones284):

so is s19=128 correct?